wiki:UC20201208

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 12/08/2020

Attending: Nancy Chanover (NMSU), Jamey Eriksen (APO), John Bally (CU), Moire Prescott (NMSU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Ben Williams (UW), Misty Bentz (GSU), Denise Stephens (BYU), Mukremin Kilic (OU), Joanne Hughes (Seattle U), Kal Kadlec (UW), Chip Kobulnicky (UWy), Derek Buzasi (FGCU), Bill Ketzeback (APO), Russet McMillan (APO)


User feedback and comments from institutional representatives

  • BYU - they had a great training session in November, got 3 half nights of good observations and got about 7 people trained (faculty and grad students) - kudos to Candace for doing such a great job with the training
  • CU - nothing to report, but John Bally had a great shift of observing this morning
  • FGCU - nothing to report, but Derek is scheduled to observe tonight so he's glad that we're open
  • Georgia State - nothing to report
  • JHU - nothing to report
  • NAPG - no report
  • NMSU - nothing to report
  • Oklahoma - nothing to report
  • Seattle - nothing to report
  • UVa - nothing to report, but the skiing this morning at Eldora was awesome :)
  • Washington - nothing to report
  • Wyoming - nothing to report

Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 11/05/2020 - 12/04/2020

0) Overview

APO was closed starting the night of 11/16 and reopened on the night of 12/2 due to the Public Health Order issued by the Governor of New Mexico.  Before 11/16 APO was operating with staffing levels based on state mandated occupancy limits as well as the NMSU return to research Phase 3 limits.  With the return on 12/2 APO is limited by the state to 25% maximum occupancy staffing levels. This 25% limit will be the case while Otero County is in the Red tier of the new Red-Yellow-Green reopening system being used in NM.  The limit of 25% is also in place for the Yellow level; it goes to 50% for the Green level.  Tiers are based on the number of cases per 100k people as well as positivity rates in each county:

Red : > 8 cases per 100k and > 5% positivity rate (both are over and average for the last 14 day period)

Yellow : < 8 cases per 100k or < 5% positivity rate (both are over and average for the last 14 day period)

Green : < 8 cases per 100k and < 5% positivity rate (both are over and average for the last 14 day period)

Here is a website that describes this system: https://cv.nmhealth.org/public-health-orders-and-executive-orders/red-to-green/

We are still restricting all visitors to the site.  Following the state of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors. Weather for the month of November has been as expected for an average year.  The last few days have been the coldest so far this year.  

1) Telescope

The telescope has been performing as expected.  Restart of operations went smoothly and science data were acquired on the first night of reopening when the weather cooperated.

2) Instruments

DIS: Regarding the scattered light, the red camera is stable and considered good to excellent; the blue camera is stable and also considered good to excellent.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The inter-order light (IOL) for red and blue continues to improve slowly with cooling weather.

NIC-FPS: The instrument was warmed up during the most recent shutdown and is on the vacuum pump now.  An estimated return to operations date is difficult to determine as the time required to reach acceptable vacuum level is the driving factor for cooling the instrument.

ARCTIC: powered on, cold, and operational. Issue with dome flats is being debugged and solutions are being worked on.  (“Fence” baffle on the mirror cover petal design is being worked on, initial design ideas had to be modified because of interference issues with the telescope structure). The new rotation redesign for the diffuser arrived on site and installation will begin when staffing levels permit.  As a reminder, the ARCTIC shutter has issues when the temp reaches -10 C.  On the first night back to operations we experienced this issue.

3) ARCSAT

ARCSAT is operational and has been used for observations this past month. Driver/software issues associated with changing temperature as well as upgraded software have settled down and appear to be resolved. Issue between the two modes of operations for pointing is being investigated.

Shelyak eShel commissioning took place at the beginning of November. Data are being analyzed and discussed from the commissioning work.


Additional telescope and instrument discussion

The telescope performance is nominal post-shutdown. We had some engineering time the night after we got back on sky so that enabled us to make some minor adjustments to the system. The temperatures got really cold and that caused some problems with the ARCTIC shutter, which fails around -10 C. Not much can be done about that for the time being. ARCSAT is operational and we have been taking this week of engineering time to dig into the issue that Muk reported on behalf of the OU users last month. It isn't solved yet but they're getting closer. John Bally reported that he learned something new with ARCES this morning, which is that you can't switch from ND5 straight to ND1 because it causes a problem in the filter wheel. The solution is to go from ND5 to <something else> and then to ND1.


APO and the COVID-19 virus

We are back to operating at a 25% on-site staffing level, which is more restrictive than what we had been doing prior to the recent shutdown in mid-November (it is the same as the operating levels when we reopened back in June after the extended spring shutdown). It may be a while before we move into the "green" zone where we can have 50% of the staff on site. Thus, the usual cautions about no on-site visitors, requesting extra patience with the site staff and timely responses to the reminder emails still apply. Nancy thanked the observatory staff for getting us back on sky and being available to help ensure the continued smooth operation of the 3.5m. We will alert the user community if there are any changes from our current mode of operations.


Q4 3.5m scheduling

There is a little bit of OPEN/DD time remaining in Q4. We have received several requests for Dec 18A and will consider all requests for that night received by the end of this week (Friday Dec 11). For all other remaining OPEN/DD time, UC reps should notify users that there is a limited amount of OPEN time Q4. If interested in requesting some time, please follow standard protocols for making such requests.


Q4 ARCSAT scheduling

The Q4 schedule is full. The only open time is the week of December 28, which is handled in the Q1 schedule. A call for Q1 ARCSAT proposals will go out next week and proposals will be due at the end of December (note the short turnaround time for anyone interested in requesting the week of December 28!).


Q1 3.5m scheduling

The Q1 2021 schedule is coming together, although it is a bit complicated. There are a couple of big monitoring programs that have requested brief blocks of time in early evening and late morning, respectively, so Russet is trying to identify programs that could benefit for slightly longer blocks that straddle the middle of the night and don't require twilight time for cals. There are not a lot of class training trips (or virtual training) requests, just one big one in March from CU. The schedule has some slight oversubscription in the early A halves but it's not too bad so far. Russet notified Bill that there is a window of time in January that could be set aside for commissioning the ARCTIC diffuser rotator.


Note Added in Proof: Updated Training Policies and Procedures

As a reminder, we are now offering one training session per quarter for PhD-level users who need to be trained for remote observing (postdocs, research staff, new faculty, etc.). Anyone interested in taking advantage of this opportunity in Q1 should notify Ben, Russet, and Candace.


ARCTIC Diffuser Rotation

The diffuser upgrade parts arrived from UW in early November. We had scheduled to commission the diffuser rotator in mid-November but then we shut down. It requires modifications of existing parts on ARCTIC and rearranging some of the electronics on the instrument. Due to the post-shutdown restrictions we have reduced availability of machine shop and engineering staff so we will likely put off commissioning the rotator until Q1. Bill hopes that it will be available to users by mid-Q1.


KOSMOS Update

The dewar seems to be holding vacuum and staying cold, which is great news. The TEG is now installing everything into it, making sure that all components are clean and that the dewar can still pump down with parts inside. So far so good. They are testing the temperature sensor and heater inside the dewar. After that they will be ready to install the detector and then begin in-lab testing. Nancy reiterated that the current plan is NOT to wait until it is safe to travel to the site for commissioning, but rather, to do as much pre-commissioning work in lab as possible before shipping the instrument and then support the on-site activities remotely. There are still many details to be worked out.


Science Communications Update

The APOLLO outreach event that was scheduled for November 19, to be live-streamed through the Las Cruces Museums Facebook page, was postponed due to the shutdown. It has been rescheduled for December 21. Nancy reported that there has not been much science promotion over the past month, partly because of the shutdown and partly because of the intern's availability. We hope to resume these activities in the near future. There was a brief discussion about the "science with APO" meeting, which will be scheduled as a virtual meeting to be held in the summer of 2021. There are many details yet to be worked out, but Nancy asked for input from those who have attended virtual conferences over the past 9 months concerning what has or has not worked well. Several suggestions were put forth, including: giving plenty of advance notice of the meeting and the deadline for signing up for talks or posters, having one person display all the presentations to reduce issues with connectivity and screen sharing, having presenters be able to share multiple applications to allow for a more in-depth kind of presentation, and the possibility of including international participants given that the meeting will be virtual. Stay tuned! This will be discussed again on a future call.


Planning for Spring Semester

We will continue meeting on the first Tuesday of the month at 10:30 am Mountain Time. Meeting dates will thus be on January 5, February 2, March 2, April 6, May 4, and June 1. Nancy will set up a repeating event in Zoom and will send out the calendar invitation.


ACTION ITEMS

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN. She just received pricing information from one outstanding request and is now ready to make a decision on which vendor to move forward with.
  • UC reps: notify users that there is a limited amount of 3.5m OPEN time in Q4. If interested in requesting some time, please follow standard protocols for making such requests.
  • Nancy: develop a survey to get feedback on the pickup training session. Status: OPEN. This is nearly done. Once it is completed, we will send it to those who were trained virtually in Q3 and Q4, as well as to the User Committee reps for distribution.

New action items from this meeting:

  • Nancy will send out the calendar invitation to the UC reps for our Jan-Jun 2021 meeting schedule.

All Other Business

Russet asked about the availability of the Shelyak on ARCSAT for Q1 2021. Bill replied that it is not quite ready yet. There are still some engineering issues to be worked out and the recent shutdown prevented some scheduled work on those. We don't want to make it available to the general user community until we have a finalized commissioning document with user instructions, performance estimates, etc.


Next meeting

The next meeting will be on Tuesday January 5 at 10:30 am MST.


Note: See TracWiki for help on using the wiki.